OpenStudy (anonymous):

if h(x)=2-x/2, find (h o h^-1)(2) A)X B)1/2 C)2 D)0 THANK YOU SO MUCH!!

jimthompson5910 (jim_thompson5910):

is the function this? \[\Large h(x) = 2 - \frac{x}{2}\] OR is the function this? \[\Large h(x) = \frac{2-x}{2}\]

OpenStudy (anonymous):

No, it is..

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

ok so the first one I wrote

OpenStudy (anonymous):

So yes, the first one you stated. Thank you.

jimthompson5910 (jim_thompson5910):

think of it as \[\Large y = 2 - \frac{x}{2}\]

jimthompson5910 (jim_thompson5910):

then swap x and y \[\Large x = 2 - \frac{y}{2}\] then solve for y

OpenStudy (anonymous):

How

OpenStudy (anonymous):

what about the h o h ^-1(2)

jimthompson5910 (jim_thompson5910):

I would first multiply both sides by 2 that clears out the fraction

jimthompson5910 (jim_thompson5910):

so you'll have 2x = 4 - y isolate y

OpenStudy (anonymous):

-y=2x-4

OpenStudy (anonymous):

then what happens with the h o h ^1 (2). That is my primary concern.

jimthompson5910 (jim_thompson5910):

-y=2x-4 will turn into y = -2x+4

jimthompson5910 (jim_thompson5910):

so the inverse is \[\Large h^{-1}(x) = -2x+4\]

jimthompson5910 (jim_thompson5910):

\[\Large (h \circ h)^{-1}(2)\] is the same as \[\Large h(h^{-1}(2))\] the first task is to compute \(\Large h^{-1}(2)\)

OpenStudy (anonymous):

Okay

jimthompson5910 (jim_thompson5910):

sorry I meant to say \(\Large (h \circ h^{-1})(2)\)

jimthompson5910 (jim_thompson5910):

anyways, what output do you get when you plug x = 2 into the inverse

OpenStudy (anonymous):

0?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Oh I see the -1

jimthompson5910 (jim_thompson5910):

then that 0 is plugged into the h(x) function

jimthompson5910 (jim_thompson5910):

\[\Large \Large h(\color{red}{h^{-1}(2)}) = h(\color{red}{0})\]

OpenStudy (anonymous):

Thank you so much for the walkthrough, it really helped me. :)

jimthompson5910 (jim_thompson5910):

so you found what h(0) is?

OpenStudy (anonymous):

0 right?

jimthompson5910 (jim_thompson5910):

plug it into 2 - (x/2)

OpenStudy (anonymous):

h^-1 (2)= -2(2)+4

jimthompson5910 (jim_thompson5910):

yeah that part is 0, but we're not done yet

OpenStudy (anonymous):

oh, so it is just 2?